OpenStudy (anonymous):

perform the indicated operation and state he domain. let f(x)=3x and g (x)=x-5 f(x)+g(x)

Directrix (directrix):

f(x)+g(x) = 3x + x - 5 = 4x - 5. Denote this as (f+g)(x) = 4x - 5.

Directrix (directrix):

The domain of a function is the set of acceptable values of x. So, @bri4life14 what would that be?

OpenStudy (anonymous):

how do you solve it

Directrix (directrix):

If the function were y = 1/x, then the domain would be the set of Real Numbers excluding zero because 1/0 is undefined. You have no denominator in your function. You have no radicals which must be kept non-negative. The question is this: Is there any value of x you should exclude from (f+g)(x)? If not, the domain is the set of Real Numbers.

OpenStudy (anonymous):

okay